OpenStudy (anonymous):

Can someone match these? 1. 45/100 = 20/x 2. 20/100 = 45/x 3. x/20 = 45/100 4. 45/20 = x/100 A. 20% of what number is 45? B. 45 is what percent of 20? C. What is 45% of 20? D. 20 is 45% of what number?

OpenStudy (anonymous):

Help ! D:

OpenStudy (lalaly):

\[A) \frac{20}{100} \times x=45 \rightarrow \frac{20}{100}=\frac{45}{x}\]which is (2)

OpenStudy (lalaly):

\[B) 45 \times \frac{x}{100}=20 \rightarrow \frac{45}{100}=\frac{20}{x}\]which is (1)

OpenStudy (lalaly):

\[C) \frac{45}{100} \times 20=x \rightarrow \frac{x}{20}=\frac{45}{100}\]which is (3)

OpenStudy (lalaly):

D with 4 :D